OpenStudy (jayden):

Solve y over -6 + 5 = 9.

OpenStudy (jackthegreatest):

so the equation is \[-6 +5 = 9?\]

OpenStudy (jackthegreatest):

or \[\frac{ y }{ -6 + 5 } = 9\]

OpenStudy (jayden):

beginning is y over -6 it wont let me type it

OpenStudy (jackthegreatest):

\[\frac{ y }{ -6 } + 5 = 9 ?\]

OpenStudy (jayden):

yes

OpenStudy (jackthegreatest):

k you would first subtract each side by 5

OpenStudy (jayden):

ok

OpenStudy (jackthegreatest):

\[\frac{ y }{ -6 } = 4\]

OpenStudy (jackthegreatest):

multiply by -6 on each side

OpenStudy (jackthegreatest):

\[y = 4 (-6)\]

OpenStudy (jackthegreatest):

y = -24

OpenStudy (jayden):

ok